OpenStudy (anonymous):

[(x-7)+6][(x-7)-6] multiply or find the special product.

OpenStudy (angela210793):

[(x-7)+6][(x-7)-6]=(x-7)^2-6^2=x^2-14x+49-36=x^2-14x+13

OpenStudy (angela210793):

(a-b)*(a+b)=a^2-b^2